# Prospects for Observing High-redshift Radio-loud Quasars in the SKA Era: Paving the Way for 21-cm Forest Observations

### v1

@@ Abstract
The 21-cm forest is a sensitive probe for the early heating process and small-scale structures during the epoch of reionization (EoR), to be realized with the upcoming Square Kilometre Array (SKA). Its detection relies on the availability of radio-bright background sources, among which the radio-loud quasars are very promising, but their abundance during the EoR is still poorly constrained due to limited observations. Here, we use a physics-driven model to forecast future radio-loud quasar observations. We fit the parameters of the model using observational data of high-redshift quasars. Assuming Eddington accretion, the model yields an average lifetime of $t_{\rm q}\sim 10^{5.3}$yr for quasars at $z\sim 6$, consistent with recent results obtained from quasar proximity zone pre-study. We show that if the radio-loud fraction of quasars evolves with redshift, it will significantly reduce the abundance of observable radio-loud quasars in the SKA era, making 21-cm forest studies challenging. With a constant radio-loud fraction, our model suggests that a one-year sky survey conducted with SKA-LOW has the capability to detect approximately 20 radio-loud quasars at $z\sim 9$, with sufficient sensitivity to resolve individual 21-cm forest lines.
@@ 1 Introduction
Probing the 21-cm signals of hydrogen atoms in the early universe mainly relies on the cosmic microwave background as the background radio source, detecting the absorption signals from the dark ages and the cosmic dawn, or the emission signals during the EoR when the gas is heated up by the early galaxies. Due to cosmic expansion, these 21-cm signals have been redshifted to low-frequency radio band, typically requiring low-frequency radio telescopes for observation. Currently, experiments to detect the cosmic dawn and EoR are in full swing, including experiments to detect the sky-averaged spectrum of the 21-cm signal during the cosmic dawn, such as Experiment to Detect the Global EoR Signature (EDGES) (Bowman et al., 2018), Shaped Antenna measurement of the background RAdio Spectrum (SARAS) (Singh et al., 2018), Hongmeng Project (Chen et al., 2023), Mapper of the IGM Spin Temperature (MIST) (Monsalve et al., 2024), The Radio Experiment for the Analysis of Cosmic Hydrogen (REACH) (de Lera Acedo et al., 2022), Probing Radio Intensity at high-Z from Marion (PRIZM) (Philip et al., 2019), The Large-aperture Experiment to Detect the Dark Age (LEDA) (Price et al., 2018), Probing ReionizATion of the Universe using Signal from Hydrogen (PRATUTH) (Sathyanarayana Rao et al., 2023), Dark Ages Polarimeter Pathfinder (DAPPR)  [^1] https://www.colorado.edu/project/dark-ages-polarimeter-pathfinder/ and the High-Z All-Sky Spectrum Experiment (HIGH-Z) (Navros, 2022). Additionally, there are experiments to detect the 21-cm power spectrum during the EoR, such as Low Frequency Array (LOFAR) (Patil et al., 2017), Murchison Widefield Array (MWA) (Tingay et al., 2013), Precision Array for Probing the Epoch of Reionization (PAPER) (Parsons et al., 2010), Hydrogen Epoch of Reionization Array (HERA) (DeBoer et al., 2017), The upgraded Global Multi-Resolution Topography (uGMRT) (Gupta et al., 2017) and the New Extension Upgrading LOFAR (NenuFAR) (Munshi et al., 2024). In particular, the Square Kilometre Array (SKA) (Huynh & Lazio, 2013) under construction will be the largest radio telescope to date. It has a large field of view, a broad band, and an ultra-high sensitivity, capable of taking tomographic images of ionizing intergalactic medium (IGM) (Koopmans et al., 2015), which will reveal more detailed information of the EoR and make an important contribution to understanding the origin of cosmic structures.
@@
In addition to these three measurement modes, another method of observing 21-cm signals is known as the 21-cm forest. If there are radio-bright point sources at the EoR, such as radio-loud quasars and radio afterglows of high-$z$ gamma-ray bursts, neutral clumps along the line of sight would create individual absorption lines on the source spectra, forming forest-like features called the 21-cm forest (in analogy to the Lyman-$\alpha$ forest) (Carilli et al., 2002; Furlanetto & Loeb, 2002; Furlanetto, 2006; Xu et al., 2009, 2010). The observation of 21-cm forest is of great value for studying the properties of the IGM, as it is highly sensitive to the heating process caused by the formation of early galaxies (Xu et al., 2009, 2011; Ciardi et al., 2013), and also for constraining the abundance of small-scale structures, so as to measure the properties of dark matter (DM) (Shimabukuro et al., 2014, 2020; Kadota et al., 2021). Recently, it was found that the one-dimensional cross power spectrum measurement of the 21-cm forest can be used to help extract faint 21-cm signals, and can also further measure the properties of DM and the early galaxies simultaneously (Shao et al., 2023).
@@
The utilization of the 21-cm forest method is presently limited by three factors. Firstly, attaining the sensitivity necessary for detecting the 21-cm forest poses a challenge. Current radio telescopes such as LOFAR and MWA, lack the capability of conducting extensive detection deep into the EoR (Ciardi et al., 2013; Chatterjee et al., 2024). The anticipated completion of SKA in the near future will significantly mitigate this existing detection challenge (Ciardi et al., 2015; Koopmans et al., 2015; Šoltinský et al., 2021). The vast effective aperture of SKA-LOW will empower us to detect the reionization process with unparalleled sensitivity. Secondly, there is a lack of an analytical model to connect physical parameters (such as dark matter mass and IGM temperature, etc.) with the 21 cm forest observational signals, making it difficult to carry out observational constraints on physical parameters (note that parameter inference based on large-scale simulations is extremely resource-intensive). To address this issue, we can consider constructing an analytical model. Alternatively, it is also possible to utilize methods based on deep learning. Recently, Sun et al. (2024) have developed a set of deep learning-based data generation and likelihood-free parameter inference methods to simultaneously address the issues of the absence of an analytical model and small-scale non-Gaussianity. Thirdly, the 21-cm forest necessitates high-redshift radio point sources as backgrounds. Owing to observational limitations, our understanding of the abundance of high-redshift radio sources remains limited, constituting the foremost uncertainty in the application of the 21-cm forest.
@@
Quasars are one of the brightest sources in the universe, with about 10% of quasars exhibiting strong radio emissions (Stern et al., 2000; Ivezić et al., 2002; Bañados et al., 2015; Gloudemans et al., 2021; Liu et al., 2021). These radio-loud quasars serve as ideal background sources for 21-cm forest. With the continuous advancement of optical and near-infrared detection, the frontier of observing quasar redshift has reached 7.64 (Wang et al., 2021). However, there are only 8 quasars observed with redshift exceeding 7 and 14 radio-loud quasars observed with redshift exceeding 6 (McGreer et al., 2006; Willott et al., 2010; Bañados et al., 2015, 2021; Ighina et al., 2021; Liu et al., 2021; Gloudemans et al., 2022; Bañados et al., 2023; Ighina et al., 2023). During the deeper reionization epochs at higher redshifts, it is crucial to obtain theoretical estimates to determine whether there are sufficient radio-loud quasars available for 21-cm forest research.
@@
To predict the abundance of high-redshift radio quasars, Haiman et al. (2004) (hereafter H04) developed a model driven by physical processes. By assuming a correspondence between supermassive black holes (SMBHs) and halos, as well as employing the Eddington limit for accretion, this model successfully derived the abundance of radio quasars consistent with observational data. After nearly two decades of equipment upgrades and observational data accumulation, we have gained further understanding of some of the physical and statistical properties of high-redshift quasars (e.g. Shang et al., 2011; Runnoe et al., 2012; Matsuoka et al., 2018; Shimasaku & Izumi, 2019; Yang et al., 2021; Matsuoka et al., 2023; Šoltinský et al., 2023; Keller et al., 2024). Currently, some assumptions of the H04 model are challenged by observational data, requiring reasonable extensions and calibrations to fit the actual distribution of high-redshift quasars. This paper aims to update this model and make reasonable predictions for radio-loud quasar observations in the SKA era.
@@
This paper is structured as follows. In Section 2, we describe the construction of the model and its calibration using high-redshift observational data. In Section 3, we present predictions for future SKA-LOW observations of radio-loud quasars and discuss the impact of currently poorly constrained parameters on the results. In Section 4, we give a summary of our findings and their implications. For this study, we have presumed the $\rm\Lambda$CDM model and have used the cosmological parameters as measured by $Planck\ \textit{2018}$ (Planck Collaboration et al., 2020): $\Omega_{m}$ = 0.3111, $\Omega_{\Lambda}$ = 0.6889, $\Omega_{b}$ = 0.03314, $h$ = 0.6766, $\sigma_{8}$ = 0.8102, and $n_{s}$ = 0.9665.
@@ 2 Method
In this section, we focus on demonstrating the construction process of the model and discussing the rationality of some physical assumptions, eventually providing a calibrated model suitable for observing high-redshift ($z\sim 6$) quasars. The calibration of the model uses quasar UV luminosity functions data at $z\sim 6$ (Matsuoka et al., 2018) and $z\sim 6.8$ (Matsuoka et al., 2023), as well as the black hole masses and the DM halo masses they inhabit obtained from 49 quasars at $z\sim 6$ (Shimasaku & Izumi, 2019).
@@ 2.1 Halo Mass Function
Quasars are powered by SMBHs. Therefore, we first need to estimate the abundance of SMBHs at the centers of galaxies. We assume that the SMBH population traces DM halos, and the distribution of DM halos follows the Sheth-Tormen form (Sheth & Tormen, 2002),
$f(\sigma,z)=0.322\sqrt{\frac{3}{2\pi}}\frac{\delta_{c}}{\sigma}\left(1+{\left(\frac{4\sigma^{2}}{3{\delta_{c}}^{2}}\right)}^{0.3}\right)\exp{\left[-\frac{3{\delta_{c}}^{2}}{8\sigma^{2}}\right]},$ (1)
where $\delta_{c}=1.6865$ is the threshold density for spherical collapse. $\sigma$ is the density fluctuation,
@@ 2.2 Halo-SMBH Mass Relation
By assuming that the radiation feedback and outflows of quasars determine the size of black holes, we can obtain the scaling relationship between black hole mass and DM halo mass (Silk & Rees, 1998; Haehnelt et al., 1998; Wyithe & Loeb, 2003). The mass scaling relationship is expressed as (Equation (4) in Wyithe & Loeb (2003)):
@@
We note that H04 used quasar luminosity function data at $z\gtrsim 3$ to fit the amplitude parameter, with the average quasar lifetime fixed at $2\times 10^{7}$ years. However, the results obtained did not match the observations of the local universe and those at redshift 6 (see Figure 1). Note that recent studies on quasar lifetimes suggest shorter lifetimes (Eilers et al., 2017, 2021; Morey et al., 2021), and in our analysis, the average quasar lifetime is treated as a variable and is constrained by high-redshift luminosity function data (see Section 2.4). To better predict the distribution of high-redshift quasars, we used the SMBH and halo masses obtained from observational data of 49 quasars at $z\sim 6$ (Shimasaku & Izumi, 2019) to fit the amplitude parameter (red solid line in Figure 1). We found that the best fit amplitude parameter $A=10^{6.83}$ is approximately 7 times the amplitude used by H04.
Figure 1: $M_{\rm halo}-M_{\rm BH}$ relation. The red lines represent $z=6$, and the blue lines represent $z=0$. Solid lines represent observational fitting, while dashed lines represent the mass relation used by H04. The shaded area represents the $1\sigma$ error. The gray triangular data points are from Ferrarese (2002), and the green data points are from Shimasaku & Izumi (2019).

### v2

@@ Abstract
The 21-cm forest is a sensitive probe for the early heating process and small-scale structures during the epoch of reionization (EoR), to be realized with the upcoming Square Kilometre Array (SKA). Its detection relies on the availability of radio-bright background sources, among which the radio-loud quasars are very promising, but their abundance during the EoR is still poorly constrained due to limited observations. In this work, we use a physics-driven model to forecast future radio-loud quasar observations. We fit the parameters of the model using observational data of high-redshift quasars. Assuming Eddington accretion, the model yields an average lifetime of $t_{\rm q}\sim 10^{5.5}$yr for quasars at $z\sim 6$, consistent with recent results obtained from quasar proximity zone observation. We show that if the radio-loud fraction of quasars evolves with redshift, it will significantly reduce the abundance of observable radio-loud quasars in the SKA era, making 21-cm forest studies challenging. With a constant radio-loud fraction, our model suggests that a one-year sky survey conducted with SKA-LOW has the capability to detect approximately 20 radio-loud quasars at $z\sim 9$, with sufficient sensitivity to resolve individual 21-cm forest lines.
@@ 1 Introduction
Probing the 21-cm signals of hydrogen atoms in the early universe mainly relies on the cosmic microwave background as the background radio source, detecting the absorption signals from the dark ages and the cosmic dawn, or the emission signals during the EoR when the gas is heated up by the early galaxies. Due to cosmic expansion, these 21-cm signals have been redshifted to low-frequency radio band, typically requiring low-frequency radio telescopes for observation. Currently, experiments to detect the cosmic dawn and EoR are in full swing, including experiments to detect the sky-averaged spectrum of the 21-cm signal during the cosmic dawn, such as the Experiment to Detect the Global EoR Signature (EDGES) (Bowman et al., 2018), the Shaped Antenna measurement of the background RAdio Spectrum (SARAS) (Singh et al., 2018), the Hongmeng Project (Chen et al., 2023), the Mapper of the IGM Spin Temperature (MIST) (Monsalve et al., 2024), the Radio Experiment for the Analysis of Cosmic Hydrogen (REACH) (de Lera Acedo et al., 2022), Probing Radio Intensity at high-Z from Marion ($\rm PRI^{Z}M$) (Philip et al., 2019), the Large-aperture Experiment to Detect the Dark Age (LEDA) (Price et al., 2018), Probing ReionizATion of the Universe using Signal from Hydrogen (PRATUTH) (Sathyanarayana Rao et al., 2023), the Dark Ages Polarimeter Pathfinder (DAPPR) [^1] https://www.colorado.edu/project/dark-ages-polarimeter-pathfinder/, and the High-Z All-Sky Spectrum Experiment (HIGH-Z) (Navros, 2022). Additionally, there are experiments to detect the 21-cm power spectrum during the EoR, such as the Low Frequency Array (LOFAR) (Patil et al., 2017), the Murchison Widefield Array (MWA) (Tingay et al., 2013), the Precision Array for Probing the Epoch of Reionization (PAPER) (Parsons et al., 2010), the Hydrogen Epoch of Reionization Array (HERA) (DeBoer et al., 2017), the upgraded Giant Metrewave Radio Telescope (uGMRT) (Gupta et al., 2017), and the New Extension in Nançay Upgrading LOFAR (NenuFAR) (Munshi et al., 2024).
In particular, the Square Kilometre Array (SKA) (Huynh & Lazio, 2013) under construction will be the largest radio telescope to date. It has a large field of view, a broad band, and an ultra-high sensitivity, capable of taking tomographic images of ionizing intergalactic medium (IGM) (Koopmans et al., 2015), which will reveal more detailed information of the EoR and make an important contribution to understanding the origin of cosmic structures.
@@
In addition to these three measurement modes, another method of observing 21-cm signals is known as the 21-cm forest. If there are radio-bright point sources at the EoR, such as radio-loud quasars and radio afterglows of high-redshift gamma-ray bursts, neutral clumps along the line of sight would create individual absorption lines on the source spectra, forming forest-like features called the 21-cm forest (in analogy to the Lyman-$\rm\alpha$ forest) (Carilli et al., 2002; Furlanetto & Loeb, 2002; Furlanetto, 2006; Xu et al., 2009, 2010). The observation of 21-cm forest is of great value for studying the properties of the IGM, as it is highly sensitive to the heating process caused by the formation of early galaxies (Xu et al., 2009, 2011; Ciardi et al., 2013), and also for constraining the abundance of small-scale structures, so as to measure the properties of dark matter (DM) (Shimabukuro et al., 2014, 2020; Kadota et al., 2021). Recently, it was found that the one-dimensional cross power spectrum measurement of the 21-cm forest can be used to help extract faint 21-cm signals, and can also further measure the properties of DM and the early galaxies simultaneously (Shao et al., 2023).
@@
The utilization of the 21-cm forest method is presently limited by three factors. Firstly, attaining the sensitivity necessary for detecting the 21-cm forest poses a challenge. Current radio telescopes such as LOFAR and MWA, lack the capability of conducting extensive detection deep into the EoR (Ciardi et al., 2013; Chatterjee et al., 2024). The anticipated completion of SKA in the near future will significantly mitigate this existing detection challenge (Ciardi et al., 2015; Koopmans et al., 2015; Šoltinský et al., 2021). The vast effective aperture of SKA-LOW will empower us to detect the reionization process with unparalleled sensitivity. Secondly, there is a lack of an analytical model to connect physical parameters (such as DM mass and IGM temperature, etc.) with the 21 cm forest observational signals, making it difficult to carry out observational constraints on physical parameters (note that parameter inference based on large-scale simulations is extremely resource-intensive). To address this issue, we can consider constructing an analytical model. Alternatively, it is also possible to utilize methods based on deep learning. Recently, Sun et al. (2024) have developed a set of deep learning-based data generation and likelihood-free parameter inference methods to simultaneously address the issues of the absence of an analytical model and small-scale non-Gaussianity. Thirdly, the 21-cm forest necessitates high-redshift radio point sources as backgrounds. Owing to observational limitations, our understanding of the abundance of high-redshift radio sources remains limited, constituting the foremost uncertainty in the application of the 21-cm forest.
@@
Quasars are one of the brightest sources in the universe, with about 10% of quasars exhibiting strong radio emissions (Stern et al., 2000; Ivezić et al., 2002; Bañados et al., 2015; Gloudemans et al., 2021; Liu et al., 2021). These radio-loud quasars serve as ideal background sources for 21-cm forest. With the continuous advancement of optical and near-infrared detection, the frontier of observing quasar redshift has reached 7.64 (Wang et al., 2021). However, there are only 9 quasars observed with redshift exceeding 7 and 14 radio-loud quasars observed with redshift exceeding 6 (McGreer et al., 2006; Willott et al., 2010; Bañados et al., 2015, 2021; Ighina et al., 2021; Liu et al., 2021; Gloudemans et al., 2022; Ighina et al., 2023; Bañados et al., 2023; Endsley et al., 2023; Banados et al., 2024). Fortunately, the successful operation of the James Webb Space Telescope (JWST) has greatly enhanced our ability to detect high-redshift active galactic nuclei (AGNs). A number of high-redshift AGNs or AGN candidates have already been reported through observations with the JWST (e.g. Übler et al., 2023; Harikane et al., 2023; Kocevski et al., 2023; Larson et al., 2023; Goulding et al., 2023; Lyu et al., 2024), making the search for AGNs at even higher redshifts possible. During the deeper reionization epochs at higher redshifts, it is crucial to obtain theoretical estimates to determine whether there are sufficient radio-loud quasars available for 21-cm forest research.
@@
To predict the abundance of high-redshift radio-loud quasars, Haiman et al. (2004) (hereafter H04) developed a model driven by physical processes. By assuming a correspondence between supermassive black holes (SMBHs) and halos, as well as employing the Eddington limit for accretion, this model successfully derived the abundance of radio-loud quasars consistent with observational data. After nearly two decades of equipment upgrades and observational data accumulation, we have gained further understanding of some of the physical and statistical properties of high-redshift quasars (e.g. Shang et al., 2011; Runnoe et al., 2012; Matsuoka et al., 2018; Shimasaku & Izumi, 2019; Yang et al., 2021; Matsuoka et al., 2023; Šoltinský et al., 2023; Keller et al., 2024). Currently, some assumptions of the H04 model are challenged by observational data, requiring reasonable extensions and calibrations to fit the actual distribution of high-redshift quasars. This paper aims to update this model and make reasonable predictions for radio-loud quasar observations in the SKA era.
@@
This paper is structured as follows. In Section 2, we describe the construction of the model and its calibration using high-redshift observational data. In Section 3, we present predictions for future SKA-LOW observations of radio-loud quasars and discuss the impact of currently poorly constrained parameters on the results. In Section 4, we give a summary of our findings and their implications. For this study, we presume the $\rm\Lambda$CDM model and use the cosmological parameters measured by Planck 2018 (Planck Collaboration et al., 2020): $\Omega_{m}$ = 0.3111, $\Omega_{\Lambda}$ = 0.6889, $\Omega_{b}$ = 0.04897, $h$ = 0.6766, $\sigma_{8}$ = 0.8102, and $n_{s}$ = 0.9665.
@@ 2 Method
In this section, we present the construction process of the model and discuss the rationality of some physical assumptions, eventually providing a calibrated model suitable for predicting high-redshift ($z\sim 6$) quasar observations. The calibration of the model uses quasar UV luminosity functions data at $z\sim 6$ (Matsuoka et al., 2018) and $z\sim 6.8$ (Matsuoka et al., 2023), as well as the black hole masses and the DM halo masses they inhabit obtained from 49 quasars at $z\sim 6$ (Shimasaku & Izumi, 2019).
@@ 2.1 Halo Mass Function
Quasars are powered by SMBHs. Therefore, we first need to estimate the abundance of SMBHs at the centers of galaxies. We assume that the SMBH population traces DM halos, and the distribution of DM halos follows the Sheth-Tormen form (Sheth & Tormen, 2002), i.e.
$\frac{{\rm d}n}{{\rm d}M_{\rm halo}}=f(\sigma,z)\ \frac{\rho_{0}}{M_{\rm halo}}\frac{{\rm d}\ {\rm ln}\left(\sigma^{-1}\right)}{{\rm d}M_{\rm halo}},$ (1)
where
$f(\sigma,z)=0.322\sqrt{\frac{3}{2\pi}}\frac{\delta_{c}}{\sigma}\left(1+{\left(\frac{4\sigma^{2}}{3{\delta_{c}}^{2}}\right)}^{0.3}\right)\exp{\left[-\frac{3{\delta_{c}}^{2}}{8\sigma^{2}}\right]}.$ (2)
Here $\delta_{c}=1.686$ is the critical overdensity for spherical collapse, and $\rho_{0}$ is the background density. $\sigma$ is the r.m.s. of density fluctuations,
@@ 2.2 Halo-SMBH Mass Relation
By assuming that the radiation feedback and outflows of quasars determine the size of black holes, we can obtain the scaling relationship between black hole mass and DM halo mass (Silk & Rees, 1998; Haehnelt et al., 1998; Wyithe & Loeb, 2003). The mass scaling relation is expressed as (Equation (4) in Wyithe & Loeb (2003)):
@@
The scaling relation between $M_{\rm BH}$ and $M_{\rm halo}$ in Equation (4) is derived froma physically motivated model combining the $M_{\rm BH}-\sigma_{\rm h}$ relation ($M_{\rm BH}\propto{\sigma_{\rm h}}^{\beta}$, where $\sigma_{\rm h}$ is halo velocity dispersion) and the halo virialization ($M_{\rm halo}\propto\sigma_{\rm h}^{3}$). The radiation-driven model proposed by Silk & Rees (1998) predicts $M_{\rm BH}\propto\sigma_{\rm h}^{5}$, while the momentum-driven model proposed by King (2003) predicts $M_{\rm BH}\propto\sigma_{\rm h}^{4}$. The results obtained from observations of the nearby universe lie between these two models (e.g. Gebhardt et al., 2000; Ferrarese & Merritt, 2000; Gültekin et al., 2009; McConnell & Ma, 2013; Kormendy & Ho, 2013), which are considered as evidence of the co-evolution of galaxies and SMBHs. However, this relation has not been thoroughly studied at high redshifts. To better predict the relation for high-redshift quasars, we use the SMBH and halo masses obtained from observational data of 49 quasars at $z\sim 6$ (Shimasaku & Izumi, 2019). We tried to treat the first power-law index in Equation (4) as a free parameter, and found that the 49 data points favor a notably low power-law index of $\sim$ 0.22. Such a low index shows large discrepancy from low-redshift observations, and it could hardly be explained by physically-motivated models. Considering the large intrinsic scatter and the large error bars of the limited sample, we attribute this discrepancy to observational systematics and incompleteness at high redshifts, which prevent us from obtaining a reliable index estimate from these data. Due to the limited understanding of the evolution of SMBHs at high redshifts, in this study we follow the choice of H04 and adopt a fixed index of 5/3.
Figure 1: $M_{\rm halo}-M_{\rm BH}$ relation. The red lines represent the fitting results at $z=6$, and the blue lines represent results for $z=0$. Solid lines represent observational fitting, while dashed lines represent the mass relation used by H04. The shaded area represents the intrinsic scatter $\sigma_{\text{in}}$. The gray triangular data points are from Ferrarese (2002), and the green data points are from Shimasaku & Izumi (2019).
We note that the $M_{\rm BH}-\sigma_{\rm bulge}$ relation measured in the nearby universe exhibits intrinsic scatter (e.g. Kormendy & Ho, 2013), which will naturally propagates to the $M_{\rm halo}-M_{\rm BH}$ relation we are fitting. Therefore, we assume that the high-redshift data contain intrinsic scatter following a Gaussian distribution $\mathcal{N}(0,\sigma_{\text{in}})$ in the logarithmic space, and construct the likelihood function as
$\mathcal{L}(A,\sigma_{\text{in}})=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi(\sigma_{m,i}^{2}+\sigma_{\text{in}}^{2})}}$
×$\times\exp\left(-\frac{\left(\ {\rm log}\left(M_{{\rm BH},i}\right)-{\rm log}\left(M_{{\rm model},i}(A)\right)\ \right)^{2}}{2(\sigma_{m,i}^{2}+\sigma_{\text{in}}^{2})}\right),$ (5)
where $A$ is the amplitude parameter defined in Equation (4), $\sigma_{\text{in}}$ is the intrinsic scatter, $\sigma_{m,i}$ is the measurement error of each data point, $M_{{\rm BH},i}$ is the black hole mass of each data point, and $M_{{\rm model},i}$ is the black hole mass obtained from Equation (4). We obtain the best-fit values of the amplitude parameter $A$ and the intrinsic scatter $\sigma_{\text{in}}$ using the maximum likelihood estimation method, and the results are shown in Figure 1 as the red line with the shaded area being the $1\sigma$ range (${\rm log}(A)=6.8\pm 0.81$), the reduced $\chi^{2}\simeq 1.02$. We found that the best-fit amplitude parameter $A=10^{6.8}$ is approximately 7 times the amplitude used by H04. As the intrinsic scatter is well-characterized by a normal distribution, the fitting result of Equation (4) represents the theoretical mean. For simplicity, we exclusively use Equation (4) with the fitted amplitude in subsequent calculations, without accounting for intrinsic scatter.
We note that H04 used quasar luminosity function data at $z\gtrsim 3$ to fit the amplitude parameter, with the average quasar lifetime fixed at $2\times 10^{7}$ years. However, the results obtained did not match the observations of the local universe and those at $z=6$ (see Figure 1). Note that recent studies on quasar lifetimes suggest shorter lifetimes (Eilers et al., 2017, 2021; Morey et al., 2021), and in our analysis, the average quasar lifetime is treated as a variable and is constrained by high-redshift luminosity function data (see Section 2.4).